if a normal human breathes at approximately 700 g of carbon dioxide per day how many O2 particles would be needed to produce that much CO2?


1
Expert's answer
2021-02-26T06:00:11-0500

Mr(CO2) -- 700 g

Mr(O2 ) -- x

x=Mr(O2 )*700/Mr(CO2)=32*700/44= 509 g

m(O2) = 509 g


n(O2)=509/32=15.9 mol

Answer: 15.9 mole of O2 would be needed to produce 700 g of carbon dioxide